14. An electrician uses 46% feet of push-back wire from a 200-foot roll. How many feet of wire are left on the roll?
Answer #1

As 46% of 200 feet is= .46×200 = 92 feet

So he uses only 92 feet wire

So wire left on roll = 200-92 = 108 feet